# Conditional Feature Importance revisited: Double Robustness, Efficiency and Inference

### v4

@@ Appendix G Inference: type-I error and power
In general, Verdinelli and Wasserman, (2024) introduced an additional square root term to ensure control of the type-I error. Indeed, due to the quadratic nature of the statistic, under the null hypothesis, the variance vanishes as $\mathbb{V}(\widehat{\psi})=O(n^{-\gamma})$ with $\gamma>1$. Therefore, to maintain a valid type-I error rate, we observe that under the null hypothesis, using Chebyshev’s inequality, we have that
@@
From Lemmas 3.11 and 3.12, we observe that in the linear setting, it is possible to use this last inequality to derive a more powerful valid test by changing the additive term $c/\sqrt{n}$ by $c/n^{-\gamma}$, with $\gamma<1$ for LOCO and $\gamma<2$ for CPI.
@@ Appendix H MSE variance estimation with influence functions
In Appendix A of Williamson et al., (2023), they propose a simple method to compute the influence function when the predictiveness measure comes from standardized V-measures. We observe that, in general, computing the empirical variance and the variance using this plug-in version do not coincide. Nevertheless, in this section, we easily prove that they do coincide when using the MSE. Indeed, we observe that
@@ Appendix I Some lemmas used in the proofs
Figure 4 illustrates that CPI converges faster under the conditional null hypothesis due to double robustness, and that a significant loss is incurred by using the data splitting method proposed in Williamson et al., (2023). However, we also observe that by increasing $n_{\mathrm{cal}}$, the double robustness of CPI can still be preserved. Additionally, the bias distribution for the non-null covariates is similar across methods. Consequently, there is improved variable selection without sacrificing variable importance.
@@
(d) Different training sample, nonnnull covariate
@@
This trade-off between variable selection and variable importance can be observed, for instance, in Figure 5. In this Figure, we study the estimated importance of two covariates in a modified version of the nonlinear setting from Bénard et al., (2022): $y=X_{0}X_{1}\mathbb{I}_{X_{2}>0}+2X_{3}X_{4}\mathbb{I}_{X_{2}<0}$, with $X\sim\mathcal{N}(\mu,\Sigma)$, where $\Sigma_{i,j}=\rho^{|i-j|}$, $p=50$, $n=10000$ and $\mu=\mathbf{0}$. The $x$-axis represents $\rho$, while the $y$-axis represents the estimated importance. The experiment is repeated 30 times.
@@
As stated in Appendix G, different variance estimators have been studied: bootstrap-based estimates for the methods with the $\mathrm{bt}$ and $\mathrm{n2}$ suffixes, and sample variance divided by the sample size for the others for the remaining methods.
When the method has no suffix (denoted by a circle marker), this indicates that no variance correction is applied. This approach was used by Chamma et al., 2024a , who ignored the vanishing variance; therefore, if the variance is zero, the null hypothesis is retained directly, without an statistical test. The suffix $\mathrm{sqrt}$ denotes a square root correction term, $\mathrm{n}$ and $\mathrm{bt}$ indicate a linear correction term, and $\mathrm{n2}$ represents a quadratic correction term.
@@
We first note that LOCO-W (Williamson et al.,, 2023) requires a larger sample size to achieve type-I error control. Consequently, while LOCO-W is capable of identifying significant covariates, it also produces a substantial number of false positives. Furthermore, we observe that Sobol-CPI(1) is the most powerful method, benefiting from its double robustness property.
@@
In Figures 9 and 10, we observe a detailed comparison of the type-I error and power, respectively. Among the methods, the most powerful tests involve the quadratic correction. Nevertheless, these do not control the type-I error. Additionally, the uncorrected method (circle marker) also fails to control the type-I error due to vanishing variance. As shown in Section J.4, the linear correction successfully preserves the type-I error.
@@
Regarding hypothesis testing, the conclusions are not entirely clear from this figure alone; we refer the reader to Figures 12 and 13 for a more detailed comparison. Nevertheless, it is evident that the Sobol-CPI methods are the most powerful, showing a clear separation in power from the LOCO methods. Moreover, the quadratic and linear corrections with bootstrap variance are not strict enough to control the type-I error. The LOCO-W method, in particular, fails to control the type-I error altogether—so even though it may yield discoveries, they are not reliable due to a high number of false positives. This can also be explained by the AUC: even with a large sample size, there is no clear distinction between important and null covariates.
@@
In Figure 12, we observe, similarly to the linear case, that the quadratic correction is not sufficient to control the type-I error. However, we also see that the linear correction with the variance estimated via bootstrap is not sufficient either. This contrasts with the linear setting, and the reason lies in the fact that the convergence rate result (see Lemma 3.11), which allows the use of Markov’s inequality (see Appendix G) to guarantee type-I error control, is only available in the linear setting.
@@
In this example we will recover the example of no-linear setting from Bénard et al., (2022) but changing the input covariance matrix to obtain more complex relationships between the covariates. Indeed, we will have $y=\alpha X^{0}X^{1}\mathbb{I}_{X^{2}>0}+\beta X^{3}X^{4}\mathbb{I}_{X^{2}<0}$, where $X$ is $p$-dimensional centered Gaussian with a Toeplitz covariance matrix where the $i,j$-th entry is given by $\rho^{|i-j|}$. In this setting, we are going to compute the $\psi_{\mathrm{TSI}}(j,P_{0})$ for the covariate $X^{0}$ and $X_{1}$.

### v5

@@ G.1 Variance correction: type-I error and power
In general, Verdinelli and Wasserman (2024) introduced an additional square root term to ensure control of the type-I error. Indeed, due to the quadratic nature of the statistic, under the null hypothesis, the variance vanishes as $\mathbb{V}(\widehat{\psi})=O(n^{-\gamma})$ with $\gamma>1$. Therefore, to maintain a valid type-I error rate, we observe that under the null hypothesis, using Chebyshev’s inequality, we have that
@@
From Lemmas 3.10 and 3.11, we observe that in the linear setting, it is possible to use this last inequality to derive a more powerful valid test by changing the additive term $c/\sqrt{n}$ by $c/n^{-\gamma}$, with $\gamma<1$ for LOCO and $\gamma<2$ for CPI.
@@
G.2 Sobol-CPI-Wilcoxon algorithm
In Algorithm 1, we provide an explicit regression-based conditional sampler to give a concrete, implementable procedure; however, any sampler could be used, and the type-I error control guarantees from Theorem 5.1 (based on the Total Variation distance) still hold.
Algorithm 1 Nonparametric Sobol-CPI (Sobol-CPI-ST/Wilcox)
0:Input: Model $\widehat{m}$, test data $\{(X_{i},y_{i})\}_{i}^{n_{\mathrm{test}}}$, control level $\alpha$ and a feature $j$
0:   (Regression) Fit $\widehat{\nu}_{j}\approx\mathbb{E}[X^{j}\mid X^{-j}]$
0:  Compute the residuals $\widehat{\epsilon}_{j}=X^{j}-\widehat{\nu}_{j}(X^{-j})$
0:  Construct
$\widetilde{X}^{\prime(j)}=\widehat{\nu}_{j}(X^{-j})+\widehat{\epsilon}_{j}^{\mathrm{permuted}}$
0:  Compute nonparametric paired test (Sign-test/Wilcoxon) between
$\left\{l\bigl(\widehat{m}(X_{i}),y_{i}\bigr)\right\}_{i}^{n_{\mathrm{test}}}\text{ and }\left\{l\bigl(\widehat{m}(\widetilde{X}^{\prime(j)}_{i}),y_{i}\bigr)\right\}_{i}^{n_{\mathrm{test}}}$
0:Return the computed p-value
G.3 Proof of Theorem 5.1 (Wilcoxon and Sign-test)
We start by recalling that the total variation for any probability measures $P_{1}$ and $P_{2}$ defined on the same probability space is given by the supremum of the difference of the measures over all measurable sets:
$d_{\mathrm{TV}}(P_{1},P_{2}):=\mathrm{sup}_{A}|P_{1}(A)-P_{2}(A)|.$
We present the proof explicitly for the Sign test for readability, but the argument for the Wilcoxon test is entirely analogous. Following Berrett et al. (2019), we observe that the total variation, conditioned on all the random variables used for the test, is exactly the distance between the estimated and original conditional distributions.
$d_{\mathrm{TV}}$ $\left\{\left(\widetilde{X}^{\prime(j)j},X_{1}^{j},\ldots,X_{n_{\mathrm{test}}}^{j})\mid\{X^{-j}_{i},y_{i}\}_{i}^{n_{\mathrm{test}}},\mathcal{D}_{\mathrm{train}}\right),\left(\widetilde{X}^{(j)j},X_{1}^{j},\ldots,X_{n_{\mathrm{test}}}^{j})\mid\{X^{-j}_{i},y_{i}\}_{i}^{n_{\mathrm{test}}},\mathcal{D}_{\mathrm{train}}\right)\right\}$
$=d_{\mathrm{TV}}\left\{\left(\widetilde{X}^{\prime(j)j})\mid\{X^{-j}_{i}\}_{i}^{n_{\mathrm{test}}},\mathcal{D}_{\mathrm{train}}\right),\left(\widetilde{X}^{(j)j})\mid\{X^{-j}_{i}\}_{i}^{n_{\mathrm{test}}},\mathcal{D}_{\mathrm{train}}\right)\right\}$
$=d_{\mathrm{TV}}\left\{P^{\prime n_{\mathrm{test}}}_{j},P^{\star n_{\mathrm{test}}}_{j}\right\},$
where we have used the assumption that the conditional sampler was estimated from a sample independent of the training sample used to fit $\widehat{m}$.
Let us define the set
$A:=\left\{\{\widetilde{X}^{(j)j}_{i}\}_{i}\in\mathcal{X}^{\times n_{\mathrm{test}}}\mid\sum l\left(\widehat{m}(\widetilde{X}^{(j)}_{i},y_{i})-l(\widehat{m}(X_{i}),y_{i})\right)\geq q^{1-\alpha}_{\mathrm{Bin}(n_{\mathrm{test}},0,5)}\right\}$
where $q^{1-\alpha}_{\mathrm{Bin}(n_{\mathrm{test}},0.5)}$ denotes the $(1-\alpha)$-quantile of a binomial distribution with parameters $n_{\mathrm{test}}$ and success probability $0.5$. Note that this corresponds to the statistic of the Sign test and can be generalized to the Wilcoxon test.
Finally, we conclude by noting that
$\mathbb{P}_{\mathcal{H}_{0}}$ $\left(\widehat{\phi}_{\mathrm{ST}}=1\mid\{X^{-j}_{i},y_{i}\}_{i}^{n_{\mathrm{test}}},\mathcal{D}_{\mathrm{train}}\right)$
$\leq P^{\star}_{j}(A)+d_{\mathrm{TV}}\left\{\left(\widetilde{X}^{\prime(j)j},X_{1}^{j},\ldots,X_{n_{\mathrm{test}}}^{j})\mid\{X^{-j}_{i},y_{i}\}_{i}^{n_{\mathrm{test}}},\mathcal{D}_{\mathrm{train}}\right),\left(\widetilde{X}^{(j)j},X_{1}^{j},\ldots,X_{n_{\mathrm{test}}}^{j})\mid\{X^{-j}_{i},y_{i}\}_{i}^{n_{\mathrm{test}}},\mathcal{D}_{\mathrm{train}}\right)\right\}$
$\leq\alpha+d_{\mathrm{TV}}\left\{P^{\prime n_{\mathrm{test}}}_{j},P^{\star n_{\mathrm{test}}}_{j}\right\}.$
@@ Appendix H MSE variance estimation with influence functions
In Appendix A of Williamson et al. (2023), they propose a simple method to compute the influence function when the predictiveness measure comes from standardized V-measures. We observe that, in general, computing the empirical variance and the variance using this plug-in version do not coincide. Nevertheless, in this section, we easily prove that they do coincide when using the MSE. Indeed, we observe that
@@ Appendix I Some lemmas used in the proofs
Figure 6 illustrates that CPI converges faster under the conditional null hypothesis due to double robustness, and that a significant loss is incurred by using the data splitting method proposed in Williamson et al. (2023). However, we also observe that by increasing $n_{\mathrm{cal}}$, the double robustness of CPI can still be preserved. Additionally, the bias distribution for the non-null covariates is similar across methods. Consequently, there is improved variable selection without sacrificing variable importance.
@@
(d) Different training sample, nonnull covariate
@@
This trade-off between variable selection and variable importance can be observed, for instance, in Figure 7. In this Figure, we study the estimated importance of two covariates in a modified version of the nonlinear setting from Bénard et al. (2022): $y=X_{0}X_{1}\mathbb{I}_{X_{2}>0}+2X_{3}X_{4}\mathbb{I}_{X_{2}<0}$, with $X\sim\mathcal{N}(\mu,\Sigma)$, where $\Sigma_{i,j}=\rho^{|i-j|}$, $p=50$, $n=10000$ and $\mu=\mathbf{0}$. The $x$-axis represents $\rho$, while the $y$-axis represents the estimated importance. The experiment is repeated 30 times.
@@
As stated in Appendix G.1, different variance estimators have been studied: bootstrap-based estimates for the methods with the $\mathrm{bt}$ and $\mathrm{n2}$ suffixes, and sample variance divided by the sample size for the others for the remaining methods.
When the method has no suffix (denoted by a circle marker), this indicates that no variance correction is applied. This approach was used by (Chamma et al., 2024a), who ignored the vanishing variance; therefore, if the variance is zero, the null hypothesis is retained directly, without an statistical test. The suffix $\mathrm{sqrt}$ denotes a square root correction term, $\mathrm{n}$ and $\mathrm{bt}$ indicate a linear correction term, and $\mathrm{n2}$ represents a quadratic correction term. $\mathrm{bt}$ uses bootstrap from the variance estimation.
@@
We first note that LOCO-W (Williamson et al., 2023) requires a larger sample size to achieve type-I error control. Consequently, while LOCO-W is capable of identifying significant covariates, it also produces a substantial number of false positives. Furthermore, we observe that Sobol-CPI(1) is the most powerful method, benefiting from its double robustness property.
@@
In Figures 11 and 12, we observe a detailed comparison of the type-I error and power, respectively. Among the methods, the most powerful tests involve the quadratic correction. Nevertheless, these do not control the type-I error. Additionally, the uncorrected method (circle marker) also fails to control the type-I error due to vanishing variance. As shown in Section J.4, the linear correction successfully preserves the type-I error.
@@
Regarding hypothesis testing, the conclusions are not entirely clear from this figure alone; we refer the reader to Figures 14 and 15 for a more detailed comparison. Nevertheless, it is evident that the Sobol-CPI methods are the most powerful, showing a clear separation in power from the LOCO methods. Moreover, the quadratic and linear corrections with bootstrap variance are not strict enough to control the type-I error. The LOCO-W method, in particular, fails to control the type-I error altogether—so even though it may yield discoveries, they are not reliable due to a high number of false positives. This can also be explained by the AUC: even with a large sample size, there is no clear distinction between important and null covariates.
@@
In Figure 14, we observe, similarly to the linear case, that the quadratic correction is not sufficient to control the type-I error. However, we also see that the linear correction with the variance estimated via bootstrap is not sufficient either. This contrasts with the linear setting, and the reason lies in the fact that the convergence rate result (see Lemma 3.10), which allows the use of Markov’s inequality (see Section G.1) to guarantee type-I error control, is only available in the linear setting.
@@
J.5 Real-data experiments
J.5.1 Datasets
For empirical evaluation, this study uses five commonly referenced benchmark datasets from the machine learning literature.
The Wisconsin Diagnostic Breast Cancer (WDBC) dataset consists of $n=569$ patient samples with $p=30$ numeric features of individual cells, obtained from a minimally invasive fine needle aspirate, to discriminate benign from malignant breast lumps. This dataset is widely used for binary classification tasks in medical informatics (Mangasarian et al., 1995).
The Wine Quality dataset (Cortez et al., 2009) comprises two related subsets: Wine-Red and Wine-White. The red wine subset contains n=1,599n=1{,}599 samples, and the white wine subset contains n=4,898n=4{,}898 samples, each with $p=11$ physicochemical input variables and a sensory quality score as the target.
The Diabetes dataset (from scikit-learn’s (Pedregosa et al., 2011) load_diabetes function) is a regression dataset with $n=442$ samples and $p=10$ variables: age, sex, body mass index, average blood pressure, and six blood serum measurements, as well as the response of interest, a quantitative measure of disease progression one year after baseline.
Finally, the California Housing dataset (Pace and Barry, 1997) contains n=20,640n=20{,}640 samples of aggregated census data with $p=8$ numeric features describing socio-economic and geographic characteristics of California districts, and a continuous target variable representing median house value.
J.5.2 Double robustness and inference
Since in real data the importance of a feature (or even knowing whether a feature is important or not) is unknown, we added an artificial feature which is a function of the original ones plus an independent noise term. Therefore, we can quantify the correlation with this feature to make the setting more complex, and we can also quantify the number of false positives by repeating the experiment across different seeds and counting the number of times this feature is selected. The discoveries are defined as the number of original features that are selected. We study the double robustness in practice, which is visible since the CFI methods assign a null importance to the artificial feature. We also study the power of the correction procedures and tests considered.
We investigate tests corrected by additive terms, which may be of order $\sqrt{n}$ as proposed by Verdinelli and Wasserman (2024), or linear as we propose for the linear setting, where the faster convergence rate can be exploited (Lemma 3.10). In particular, we also tested quadratic corrections, which failed to control the error. Additionally, we considered nonparametric tests, namely the Wilcoxon and the Sign test. We ran these simulations over all datasets from Appendix J.5.1 and across standard ML models: Lasso, Gradient Boosting, Random Forest, Neural Network, and a SuperLearner combining the previous ones. Since the results were consistent across all experiments, we only report those for the SuperLearner on WDBC (Figure 16) and red wine quality (Figure 17).
In Figure 16, we first observe that, due to double robustness, methods that do not rely on refitting are able to detect the null feature faster and with less variance. We also observe that Sobol-CPI(1) with the Wilcoxon test is the most powerful test for any correlation level, while methods based on additive corrections are too conservative. We further observe that the raw methods based on a t-test (Sobol-CPI(1) or LOCO) fail to control the error due to vanishing variance. Finally, we point out the clear time gain obtained by avoiding refitting a SuperLearner for each feature.
Figure 16: Double robustness and inference on the WDBC dataset with SuperLearner: Sobol-CPI-based methods do not assign any importance to the null artificial feature, and Sobol-CPI(1) combined with the Wilcoxon test is both the most powerful and the fastest thanks to avoiding the refitting of the SuperLearner.
Similar conclusions can be drawn for the red wine dataset (Figure 17). Indeed, it is clear not only that the Sobol-CPI methods do not assign any importance to the artificial feature, but also that they are, in general, much more powerful. Moreover, the Wilcoxon test applied to Sobol-CPI(1) is the most powerful.
Figure 17: Double robustness and inference on the Red Wine dataset with SuperLearner: Sobol-CPI-based methods do not assign any importance to the null artificial feature, and Sobol-CPI(1) combined with the Wilcoxon test is both the most powerful and the fastest.
J.5.3 Double robustness and Asymptotic relevance
In Figure 18 we report the importance of the additional null feature with respect to its correlation with the original features, across all real datasets from Appendix J.5.1 and standard ML models (Lasso, Gradient Boosting, Random Forest, Neural Network, and their SuperLearner). We evaluate Sobol-CPI(1/10/100), the standard LOCO, its data-splitting variant LOCO-W from (Williamson et al., 2023), and the Gaussian conditional samplers from fippy producing CFI and scSAGEvf, which correspond to Sobol-CPI(1) and Sobol-CPI(50) respectively, without the bias correction from Lemma 3.10, which we manually add.
The double robustness property persists: Sobol–CPI-based methods exhibit a markedly smaller variance than refitting-based LOCO methods and assign (almost) zero importance to the null feature across all datasets and correlation levels, for both regression and Gaussian conditional samplers.
We also study PFI in connection with the asymptotic relevance assumption. PFI typically assigns no importance to the null feature (as visible from the scale), except for high correlations under Neural Networks. This indicates that while the model does not utilize the null feature, there can still be nontrivial gains from conditional samplers. In addition, we recall that PFI does not produce valid $p$-values, making it unsuitable for inference. In contrast, for the CFI we have provided valid inference procedures, such as the Wilcoxon test and the variance correction, enabling hypothesis testing for variable importance.
Figure 18: Double robustness and asymptotic relevance in real data. Across all real datasets and models, Sobol-CPI methods assign zero importance to the artificial null feature for every correlation level. This also holds for other sampling schemes such as the Gaussian one (CFI $\equiv$ Sobol-CPI(1) and scSAGEvf $\equiv$ Sobol-CPI(50)), illustrating the double robustness property. In contrast, PFI, which can be viewed as a CFI with a severely misspecified imputer, still yields small importance on this null feature, except for high correlations under neural networks, consistent with the asymptotic relevance assumption.
@@
In this example we will recover the example of no-linear setting from Bénard et al. (2022) but changing the input covariance matrix to obtain more complex relationships between the covariates. Indeed, we will have $y=\alpha X^{0}X^{1}\mathbb{I}_{X^{2}>0}+\beta X^{3}X^{4}\mathbb{I}_{X^{2}<0}$, where $X$ is $p$-dimensional centered Gaussian with a Toeplitz covariance matrix where the $i,j$-th entry is given by $\rho^{|i-j|}$. In this setting, we are going to compute the $\psi_{\mathrm{TSI}}(j,P_{0})$ for the covariate $X^{0}$ and $X_{1}$.